# Using Open Source EDA Tools in ASICs for HEP: A Mixed Comparison

### v2

@@ 1 Introduction
Application-specific integrated circuits (ASICs) are central to front-end and data-acquisition electronics in high-energy-physics (HEP) experiments, where tight constraints on radiation tolerance, power, area, and latency motivate custom designs [LHCbUpgrade2024, CMSRun3_2024, HGCAL_Vertical_2024]. At the same time, design costs and license availability can limit iteration speed and collaboration. Open-source electronic design automation (EDA) flows and open process design kits (PDKs) aim to lower these barriers and improve transparency, with manufacturable demonstrations in recent tapeouts [OpenROAD_DAC19, OpenLane_WOSET20, SkyWaterPDK, IHP_OpenPDK].
Despite progress in open digital implementation, published side-by-side comparisons with commercial environments in mature technologies relevant to HEP remain limited, especially for quality of results in standard-cell logic and for the feasibility of analog layout. Existing comparative studies focus almost exclusively on digital logic, typically use earlier generations of open-source flows, and do not address analog layout or HEP-specific constraints (e.g. physical-design benchmarks in qflow, RISC-V implementations comparing OpenLane to commercial tools, and FIFO cores evaluated in qflow versus Cadence Encounter) [Sangani2022, Hesham2021, Acharya2022]. Analog physical design remains challenging; recent academic tools show promise but require careful handling of device matching and symmetry [MAGICAL_ICCAD19, ALIGN_arXiv20].
@@ 2 Methods and Flows
All designs target the IHP SG13G2 open PDK with consistent process–voltage–temperature (PVT) corners, libraries, and I/O assumptions. Fairness is enforced with identical register-transfer level descriptions or schematics and a shared Synopsys Design Constraints (SDC) file. We report standard digital quality of results (post-route area, power, timing slack, maximum frequency, cell count, congestion), wall-clock runtime, and the number of manual iterations to timing closure, where a manual iteration is a full place-and-route rerun after edits to constraints, scripts, or floorplan. For the VCO, analog layout effort is given qualitatively in person-hours of constraint-aware placement, routing, and parasitic tuning. Measured data accompany simulations when available. Tool and PDK versions are recorded for reproducibility, following open digital-flow practice [OpenROAD_DAC19, RePlAce_TCAD2018, TritonRoute_TCAD2021, TritonRouteWXL_TCAD2022].
@@ 2.1 Digital Flow
The open-source pipeline uses Yosys+ABC for logic synthesis and OpenLane/OpenROAD for place-and-route and optimization [OpenROAD_DAC19, RePlAce_TCAD2018, TritonRoute_TCAD2021, TritonRouteWXL_TCAD2022]. Static timing analysis and power estimation use SAIF or VCD activity from gate-level simulation under the IEEE 1800 waveform standard [OpenTimer_v2_IEEE_DandT_2021, ieee1800_2023], and the reported digital power figures rely on this vector-based estimation under identical stimuli in both flows. Physical verification uses the PDK KLayout decks and Netgen for LVS; the commercial baseline mirrors these stages with Genus, Innovus, and Tempus. Both flows share the same 50 MHz SDC with 0.2 ns of skew and 0.3 ns of transition time, it also includes routing-layer limits, antenna repair, and matched floorplan and utilization. Vector-based power is preferred; any vectorless estimates are explicitly identified. Runs use fixed seeds where available and a uniform runtime script, and timing-accurate SDF simulation was available only in the commercial environment.
@@ 2.2 Analog Flow
Schematic capture and pre-layout simulation use equivalent environments with identical models, stimuli, and PVT corners (Xschem+Ngspice and Virtuoso+Spectre). The benchmark is a differential current-starved ring-oscillator VCO. Oscillation frequency is taken from steady-state transient segments after startup, $K_{\mathrm{VCO}}$ from the slope of $f_{\mathrm{osc}}(V_{\mathrm{CTRL}})$ near nominal bias, and supply current from the same runs with matched numerical controls. Layout applies the same constraint set in both ecosystems: symmetric and matched devices, common-centroid or interdigitated arrangements where appropriate, and consistent use of guard rings, dummies, shielded differential nets, preferred-direction routing, and local decoupling, reflecting current evidence and challenges in automated analog physical design [MAGICAL_SiliconProven_2022, ALIGN_DAC2019, ALIGN_ASPDAC2023, ASPDAC2024_PerfDrivenAnalogPD]. Parasitic extraction uses equivalent settings (Magic ext/ext2spice or KLayout-PEX on the open side and vendor LPE on the commercial side) with back-annotated post-layout simulation. To limit bias, we equalize inputs, constraints, PVT corners, floorplan and utilization, routing-layer limits, and power activity across flows, document tool-specific differences, and repeat runs on the same host with fixed seeds when possible, and present measured and simulated data side by side.
@@ 3.1 Common-Mode Noise (CMN) Filter
The CMN filter is a digital pre-processor that estimates the common-mode component across $N$ parallel channels and subtracts it from each channel. A hardware median-finding strategy makes the estimate robust to outliers (figure 1a); a rank-based estimator selects the sample of order $\lceil N/2\rceil$. The block was developed within the SALSA front-end readout ASIC [Salsa] under a 50 MHz pipeline clock and performance, power, and area (PPA) constraints, which motivated a fully combinatorial median solution [filterscomp]. We use a Combinatorial Sum Median Finder (CSMF) that compares channel pairs in parallel and accumulates per-channel Hamming weights to select the median (figure 1b), yielding a compact combinatorial benchmark for synthesis and place-and-route.
@@ 3.2 Finite-State Machine (FSM)
The packetizer FSM orchestrates the SALSA output stream [Salsa]. A six-state controller emits frames with header, subheader, data words, and two trailer words (figure 1c), while a shift register snapshots internal registers and formats 32-bit words with timing, identification, and integrity fields. Designed under a 50 MHz clock and PPA constraints, this sequential workload stresses register placement, clock-tree synthesis with timing targets, and hold fixing on short paths, complementing the CMN benchmark.

### v3

@@ 1 Introduction
Application-specific integrated circuits (ASICs) are central to front-end and data-acquisition electronics in high-energy-physics (HEP) experiments, where tight constraints on radiation tolerance, power, area, and latency motivate custom designs [1, 9, 20]. At the same time, design costs and license availability can limit iteration speed and collaboration. Open-source electronic design automation (EDA) flows and open process design kits (PDKs) aim to lower these barriers and improve transparency, with manufacturable demonstrations in recent tapeouts [3, 6, 7, 13].
Despite progress in open digital implementation, published side-by-side comparisons with commercial environments in mature technologies relevant to HEP remain limited, especially for quality of results in standard-cell logic and for the feasibility of analog layout. Existing comparative studies focus almost exclusively on digital logic, typically use earlier generations of open-source flows, and do not address analog layout or HEP-specific constraints (e.g. physical-design benchmarks in qflow, RISC-V implementations comparing OpenLane to commercial tools, and FIFO cores evaluated in qflow versus Cadence Encounter) [18, 10, 2]. Analog physical design remains challenging; recent academic tools show promise but require careful handling of device matching and symmetry [21, 5].
@@ 2 Methods and Flows
All designs target the IHP SG13G2 open PDK with consistent process–voltage–temperature (PVT) corners, libraries, and I/O assumptions. Fairness is enforced with identical register-transfer level descriptions or schematics and a shared Synopsys Design Constraints (SDC) file. We report standard digital quality of results (post-route area, power, timing slack, maximum frequency, cell count, congestion), wall-clock runtime, and the number of manual iterations to timing closure, where a manual iteration is a full place-and-route rerun after edits to constraints, scripts, or floorplan. For the VCO, analog layout effort is given qualitatively in person-hours of constraint-aware placement, routing, and parasitic tuning. Measured data accompany simulations when available. Tool and PDK versions are recorded for reproducibility, following open digital-flow practice [3, 4, 14, 15].
@@ 2.1 Digital Flow
The open-source pipeline uses Yosys+ABC for logic synthesis and OpenLane/OpenROAD for place-and-route and optimization [3, 4, 14, 15]. Static timing analysis and power estimation use SAIF or VCD activity from gate-level simulation under the IEEE 1800 waveform standard [11, 12], and the reported digital power figures rely on this vector-based estimation under identical stimuli in both flows. Physical verification uses the PDK KLayout decks and Netgen for LVS; the commercial baseline mirrors these stages with Genus, Innovus, and Tempus. Both flows share the same 50 MHz SDC with 0.2 ns of skew and 0.3 ns of transition time, it also includes routing-layer limits, antenna repair, and matched floorplan and utilization. Vector-based power is preferred; any vectorless estimates are explicitly identified. Runs use fixed seeds where available and a uniform runtime script, and timing-accurate SDF simulation was available only in the commercial environment.
@@ 2.2 Analog Flow
Schematic capture and pre-layout simulation use equivalent environments with identical models, stimuli, and PVT corners (Xschem+Ngspice and Virtuoso+Spectre). The benchmark is a differential current-starved ring-oscillator VCO. Oscillation frequency is taken from steady-state transient segments after startup, $K_{\mathrm{VCO}}$ from the slope of $f_{\mathrm{osc}}(V_{\mathrm{CTRL}})$ near nominal bias, and supply current from the same runs with matched numerical controls. Layout applies the same constraint set in both ecosystems: symmetric and matched devices, common-centroid or interdigitated arrangements where appropriate, and consistent use of guard rings, dummies, shielded differential nets, preferred-direction routing, and local decoupling, reflecting current evidence and challenges in automated analog physical design [23, 16, 19, 22]. Parasitic extraction uses equivalent settings (Magic ext/ext2spice or KLayout-PEX on the open side and vendor LPE on the commercial side) with back-annotated post-layout simulation. To limit bias, we equalize inputs, constraints, PVT corners, floorplan and utilization, routing-layer limits, and power activity across flows, document tool-specific differences, and repeat runs on the same host with fixed seeds when possible, and present measured and simulated data side by side.
@@ 3.1 Common-Mode Noise (CMN) Filter
The CMN filter is a digital pre-processor that estimates the common-mode component across $N$ parallel channels and subtracts it from each channel. A hardware median-finding strategy makes the estimate robust to outliers (figure 1a); a rank-based estimator selects the sample of order $\lceil N/2\rceil$. The block was developed within the SALSA front-end readout ASIC [17] under a 50 MHz pipeline clock and performance, power, and area (PPA) constraints, which motivated a fully combinatorial median solution [8]. We use a Combinatorial Sum Median Finder (CSMF) that compares channel pairs in parallel and accumulates per-channel Hamming weights to select the median (figure 1b), yielding a compact combinatorial benchmark for synthesis and place-and-route.
@@ 3.2 Finite-State Machine (FSM)
The packetizer FSM orchestrates the SALSA output stream [17]. A six-state controller emits frames with header, subheader, data words, and two trailer words (figure 1c), while a shift register snapshots internal registers and formats 32-bit words with timing, identification, and integrity fields. Designed under a 50 MHz clock and PPA constraints, this sequential workload stresses register placement, clock-tree synthesis with timing targets, and hold fixing on short paths, complementing the CMN benchmark.